For the following reaction, 4.82 grams of chlorine gas are mixed with excess bromine . The reaction yields 10.4 grams of bromine monochloride .

What is the theoretical yield of bromine monochloride ?

What is the percent yield for this reaction ?
1
Expert's answer
2019-08-01T03:12:57-0400

Cl2 + Br2 = 2 BrCl


moles of chlorine gas:

"n = \frac {m}{MW} = \frac {4.82 g}{71.0 g/mole} =0.0679 moles"

moles of bromine monochloride:


"0.0679 \cdot \frac {2 mole} {1 mole}=0.136 moles"


theoretical mass of bromine monochloride:


"m= n \cdot MW = 0.136 \cdot 115.5 =15.7 g"


theoretical yield:


"E= \frac {m (practical)} {m (theoretical)} \cdot 100 \% = \frac {10.4 g} {15.7} \cdot 100 \% = 66.3 \%"